# CAMEF: Causal-Augmented Multi-Modality Event-Driven Financial Forecasting by Integrating Time Series Patterns and Salient Macroeconomic Announcements

### v1

@@ Abstract.
Accurately forecasting the impact of macroeconomic events is critical for investors and policymakers. Salient events like monetary policy decisions and employment reports often trigger market movements by shaping expectations of economic growth and risk, thereby establishing causal relationships between events and market behavior. Existing forecasting methods typically focus either on textual analysis or time-series modeling, but fail to capture the multi-modal nature of financial markets and the causal relationship between events and price movements. To address these gaps, we propose CAMEF (Causal-Augmented Multi-Modality Event-Driven Financial Forecasting), a multi-modality framework that effectively integrates textual and time-series data with a causal learning mechanism and an LLM-based counterfactual event augmentation technique for causal-enhanced financial forecasting. Our contributions include: (1) a multi-modal framework that captures causal relationships between policy texts and historical price data; (2) a new financial dataset with six types of macroeconomic releases from 2008 to April 2024, and high-frequency real trading data for five key U.S. financial assets; and (3) an LLM-based counterfactual event augmentation strategy. We compare CAMEF to state-of-the-art transformer-based time-series and multi-modal baselines, and perform ablation studies to validate the effectiveness of the causal learning mechanism and event types.  [^1] Open-Source Promise: We plan to release the CAMEF dataset and model to the research community to promote reproducibility and stimulate future research in this interdisciplinary domain upon publication.
Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Economicsccs: Computing methodologies Neural networks
@@ 1. Introduction
Previous studies on event-driven forecasting have primarily adopted three lines of methodologies. The first line of approaches utilizes text feature-based models, where language models, ranging from self-crafted RNN-based architectures (Huynh et al., 2017; Liu, 2018a; Daigo Tashiro and Sakaji, 2019; Xu and Cohen, 2018) to pre-trained transformers (Zhou et al., 2021a; Shah et al., 2023), embed sentiment information into text vectors, and then stock movements are predicted as a binary classification task (e.g., hawkish vs. dovish). The second line of methodology focuses on historical time-series data, treating stock price movements as a regression problem (Taylor, 2007; Cao and Tsay, 1992). Recently, transformer-based architectures have been applied for time-series prediction, including Informer (Zhou et al., 2021b), FedFormer (Zhou et al., 2022), and AutoFormer (Chen et al., 2021), etc. However, both of these directions typically focus on a single modality, neglecting multi-dimensional information. The third line of research adopts a multi-modality approach, leveraging multiple types of data sources to enhance forecasting performance. For instance, studies like (Mathur et al., 2022; Ouyang et al., 2024) incorporate textual, video, and audio data from FOMC meetings alongside corresponding market movements. While these approaches show promise for event-driven financial forecasts, they face three major limitations:
  • Data Limitation: Existing approaches predominantly focus on a single type of event, such as FOMC meetings (Shah et al., 2023; Mathur et al., 2022; Ouyang et al., 2024), while neglecting other crucial macroeconomic events like unemployment insurance releases, CPI, PPI, and GDP advance reports. Additionally, many studies rely on daily-based time-series data for financial assets (Shah et al., 2023; Mathur et al., 2022; Ouyang et al., 2024; Taylor, 2007; Cao and Tsay, 1992; Zhou et al., 2021b, 2022; Chen et al., 2021), which limits their applicability and precision in real-time trading scenarios where high-frequency data is mostly adopted.
  • Modality Limitation: Most prior studies rely on single-modality analysis, using either textual models (Huynh et al., 2017; Liu, 2018a; Daigo Tashiro and Sakaji, 2019; Xu and Cohen, 2018; Zhou et al., 2021a; Shah et al., 2023) or time-series models (Taylor, 2007; Cao and Tsay, 1992; Zhou et al., 2021b, 2022; Chen et al., 2021), which fail to integrate the complementary strengths of both modalities. While some multi-modality approaches have been proposed (Mathur et al., 2022; Ouyang et al., 2024), they often lack advanced mechanisms for feature fusion, effective decoding strategies, and causal learning, which are critical for understanding the complex interplay between event texts and market dynamics.
@@ 2.1. Event-Drive Financial Forecasting
Event-driven financial forecasting (Bao et al., 2025) focuses on predicting asset prices (Goldberg and Grisse, 2013; Gilbert et al., 2010) and market volatility (LUCCA and MOENCH, 2015; CIESLAK et al., 2019) based on events like macroeconomic releases (Gilbert et al., 2010), news (Liu, 2018b), corporate announcements (Zhou et al., 2021a), and social media activity (Xu and Cohen, 2018). Three main approaches exist in this area. The first leverages text analysis to predict asset responses based on event-related text. Early works utilized TF-IDF (Pui Cheong Fung et al., 2003; Li et al., 2014) and topic models (Si et al., 2013; Nguyen et al., 2015), progressing to RNN-based models (Huynh et al., 2017; Liu, 2018a) and pre-trained transformers (Zhou et al., 2021a; Shah et al., 2023), which capture nuanced semantics. Although these models excel at semantic extraction, they often lack integration with historical price data, crucial for holistic forecasting.
The second line of approaches uses statistical and sequential models on numerical data, such as linear regression (Bhuriya et al., 2017), ARIMA (Ariyo et al., 2014), and GARCH (Hyup Roh, 2007). Later, deep learning methods like RNNs (Liu, 2018a) and CNNs (Selvin et al., 2017; Durairaj and Mohan, 2022) enhanced nonlinear modeling capabilities. More recently, transformer-based models, such as Informer (Zhou et al., 2021b) and FedFormer (Zhou et al., 2022), improved long-range dependency modeling for time series data. However, these models tend to be “case-specific,” requiring task-specific training. In contrast, the lastest pre-trained models for time-series data, like MOMENT (Goswami et al., 2024), Timer (Liu et al., 2024b), and TOKEN (Anonymous, 2024), offer more generalized and adaptable solutions for time-series tasks.
@@ A. Macroeconomic Event Summary
FOMC Html Quarterly 1993.3 $\sim$2024.6 255 2,550 www.federal reserve.gov
Unemployment Insurance Claims PDF, Txt Weekly 2002.10 $\sim$2024.6 913 9,130 oui.doleta.gov
@@ B. Time-Series Data Summary
US Treasury Bond at 1-Month (USGG1M) Open, Close, High, Low 5 Min 2013.01 $\sim$2024.06 751,443
US Treasury Bond at 5-Year (USGG5YR) Open, Close, High, Low 5 Min 2013.01 $\sim$2024.06 734,773
@@ 3. Problem Formulation
Time Series Data is defined as $\mathcal{X}:=\{X_{1},X_{2},\dots,X_{|\mathcal{X}|}\}$, where each $X_{i}$ represents the numerical data at time step $i$. An event $\mathcal{E}_{i}$ is aligned with a time series segment $\mathcal{X}_{i-\tau:i+\tau}$, where $i$ denotes the time of the releasement of the event, and $\tau$ represents the duration of the time-series segment both preceding and succeeding time step $i$, denoted as $[\mathcal{X}_{i-\tau:i+\tau}\mapsto\mathcal{E}_{i}]$. This alignment reflects the time series segment leading up to the event ($\mathcal{X}_{i-\tau}$) and the period during which the event is expected to have an effect ($\mathcal{X}_{i+\tau}$).
@@ 4.1. Dataset Acquisition
In addition to the textual data, studies (CIESLAK et al., 2019; Gertler and Horvath, 2018; Masciandaro et al., 2024; Rosa, 2016; Tadle, 2022; Rosa, 2013; LUCCA and MOENCH, 2015; Rosa, 2011) have demonstrated that the largest market impacts are typically observed in major U.S. stock indexes and Treasury bonds. Therefore, we focused on collecting high-frequency trading time-series data at 5 minute interval for key stock indexes, including the S&P 500 (SPX), Dow Industrial (INDU), NASDAQ (NDX), as well as U.S. Treasury Bond at 1-Month (USGG1M) and Treasury Bond at 5-Year (USGG5YR) from Bloomberg Terminal.
@@ 4.2. Counterfactual Events Generation based on LLM
5. CAMEP Architecture
The CAMEP model integrates both textual and time-series information through a structured architecture consisting of a textual encoder, a time-series encoder, and a forecasting decoder, as depicted in Fig. 3. Each component is detailed below.
@@ 5.1. Textual Modality Encoder ($\textbf{CAMEF}_{\text{Textual}}$)
We encode event scripts using RoBERTa (Liu et al., 2020). Given an input script $\mathbf{E}_{i}={w_{1},w_{2},\dots,w_{m}}$, RoBERTa produces contextual token embeddings:
@@
where $\mathbf{W}^{(1)}\in\mathbb{R}^{768\times 1024}$, $\mathbf{W}^{(2)}\in\mathbb{R}^{1024\times 1024}$, and $\mathbf{W}^{(3)}\in\mathbb{R}^{1024\times 1024}$. The final encoding vector $\mathbf{E}_{i}$ for the script is computed as the average of the transformed embeddings, $\mathbf{E}_{i}=\frac{1}{m}\sum_{j=1}^{m}\mathbf{e}_{j}$.
@@ 5.2. Time-Series Modality Encoder
where $\mathbf{W}_{1},\mathbf{W}_{2},\mathbf{W}_{3}\in\mathbb{R}^{1024\times 1024}$ are the weight matrices, and $\mathbf{b}_{1},\mathbf{b}_{2},\mathbf{b}_{3}\in\mathbb{R}^{1024}$ are the respective biases. Finally, $\mathbf{Z}_{i}\in\mathbb{R}^{1024}$ serves as the refined vector for the time series segment.
@@ 5.3. Feature Fusion and Time Series Decoder
After obtaining the encoded vectors from both the textual data and the time series data, denoted as $\mathbf{E_{i}}$ and $\mathbf{Z}_{i}$, respectively, we concatenate these two vectors to create a unified representation:
(7) $\mathbf{E}_{combined}=\textbf{concat}(\mathbf{E_{i}},\mathbf{Z}_{i}),$
where $\mathbf{E}_{combined}$ captures both the semantic information from the macroeconomic releases and the temporal patterns from the time series data. We employ GPT-2 (Radford et al., 2019) as the decoder to decode the fused vector by leveraging its effective auto-regressive ability:
@@
where $\mathbf{H}^{(0)}=\mathbf{E}_{\text{combined}}$, and $f_{\text{GPT2\_layer}}^{(l)}$ represents the transformation function of the $l$-th GPT-2 layer, where $l=12$. After the final layer, the output is normalized using a layer normalization function:
@@ 5.4. Time-Series Forecasting Post-Regressor and Learning Objectives
[…]
@@
Model / Datasets Forecasting Length SP500 (SPX) Dow Industrial (INDU) NASDAQ (NDX) USGG1M USGG5YR
@@ 6.1.3. Test Settings
Textual Causal Feature Fusion GPT2 Decoder Post- Regressor 35 70 140 35 70 140 35 70 140 35 70 140 35 70 140 35 70 140
@@ 6.1.4. Implementation Overview
6.2. Experimental Results for Event-Driven Time-Series Forecasting (Q1)
@@
6.3. Ablation Studies on Model Components (Q2)

### v2

@@ Abstract.
Accurately forecasting the impact of macroeconomic events is critical for investors and policymakers. Salient events like monetary policy decisions and employment reports often trigger market movements by shaping expectations of economic growth and risk, thereby establishing causal relationships between events and market behavior. Existing forecasting methods typically focus either on textual analysis or time-series modeling, but fail to capture the multi-modal nature of financial markets and the causal relationship between events and price movements. To address these gaps, we propose CAMEF (Causal-Augmented Multi-Modality Event-Driven Financial Forecasting), a multi-modality framework that effectively integrates textual and time-series data with a causal learning mechanism and an LLM-based counterfactual event augmentation technique for causal-enhanced financial forecasting. Our contributions include: (1) a multi-modal framework that captures causal relationships between policy texts and historical price data; (2) a new financial dataset with six types of macroeconomic releases from 2008 to April 2024, and high-frequency real trading data for five key U.S. financial assets; and (3) an LLM-based counterfactual event augmentation strategy. We compare CAMEF to state-of-the-art transformer-based time-series and multi-modal baselines, and perform ablation studies to validate the effectiveness of the causal learning mechanism and event types.
Multimodal learning, Causal Learning, Financial dataset, Time-series Forecasting
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736872isbn: 979-8-4007-1454-2/2025/08ccs: Applied computing Economicsccs: Computing methodologies Neural networks
@@ 1. Introduction
Previous studies on event-driven forecasting have primarily adopted three lines of methodologies. The first line of approaches utilizes text feature-based models, where language models, ranging from self-crafted RNN-based architectures (Huynh et al., 2017; Liu, 2018; Daigo Tashiro and Sakaji, 2019; Xu and Cohen, 2018) to pre-trained transformers (Zhou et al., [n. d.]; Shah et al., 2023), embed sentiment information into text vectors, and then stock movements are predicted as a binary classification task (e.g., hawkish vs. dovish). The second line of methodology focuses on historical time-series data, treating stock price movements as a regression problem (Taylor, 2007; Cao and Tsay, 1992). Recently, transformer-based architectures have been applied for time-series prediction, including Informer (Zhou et al., 2021), FedFormer (Zhou et al., 2022), and AutoFormer (Chen et al., 2021), etc. However, both of these directions typically focus on a single modality, neglecting multi-dimensional information. The third line of research adopts a multi-modality approach, leveraging multiple types of data sources to enhance forecasting performance. For instance, studies like (Mathur et al., 2022; Ouyang et al., 2024) incorporate textual, video, and audio data from FOMC meetings alongside corresponding market movements. While these approaches show promise for event-driven financial forecasts, they face three major limitations:
  • Data Limitation: Existing approaches predominantly focus on a single type of event, such as FOMC meetings (Shah et al., 2023; Mathur et al., 2022; Ouyang et al., 2024), while neglecting other crucial macroeconomic events like unemployment insurance releases, CPI, PPI, and GDP advance reports. Additionally, many studies rely on daily-based time-series data for financial assets (Shah et al., 2023; Mathur et al., 2022; Ouyang et al., 2024; Taylor, 2007; Cao and Tsay, 1992; Zhou et al., 2021, 2022; Chen et al., 2021), which limits their applicability and precision in real-time trading scenarios where high-frequency data is mostly adopted.
  • Modality Limitation: Most prior studies rely on single-modality analysis, using either textual models (Huynh et al., 2017; Liu, 2018; Daigo Tashiro and Sakaji, 2019; Xu and Cohen, 2018; Zhou et al., [n. d.]; Shah et al., 2023) or time-series models (Taylor, 2007; Cao and Tsay, 1992; Zhou et al., 2021, 2022; Chen et al., 2021), which fail to integrate the complementary strengths of both modalities. While some multi-modality approaches have been proposed (Mathur et al., 2022; Ouyang et al., 2024), they often lack advanced mechanisms for feature fusion, effective decoding strategies, and causal learning, which are critical for understanding the complex interplay between event texts and market dynamics.
@@ 2.1. Event-Drive Financial Forecasting
Event-driven financial forecasting (Bao et al., 2025) focuses on predicting asset prices (Goldberg and Grisse, 2013; Gilbert et al., 2010) and market volatility (LUCCA and MOENCH, 2015; CIESLAK et al., 2019) based on events like macroeconomic releases (Gilbert et al., 2010), news (Liu, 2018), corporate announcements (Zhou et al., [n. d.]), and social media activity (Xu and Cohen, 2018). Three main approaches exist in this area. The first leverages text analysis to predict asset responses based on event-related text. Early works utilized TF-IDF (Pui Cheong Fung et al., 2003; Li et al., 2014) and topic models (Si et al., 2013; Nguyen et al., 2015), progressing to RNN-based models (Huynh et al., 2017; Liu, 2018) and pre-trained transformers (Zhou et al., [n. d.]; Shah et al., 2023), which capture nuanced semantics. Although these models excel at semantic extraction, they often lack integration with historical price data, crucial for holistic forecasting.
The second line of approaches uses statistical and sequential models on numerical data, such as linear regression (Bhuriya et al., 2017), ARIMA (Ariyo et al., [n. d.]), and GARCH (Hyup Roh, 2007). Later, deep learning methods like RNNs (Liu, 2018) and CNNs (Selvin et al., 2017; Durairaj and Mohan, 2022) enhanced nonlinear modeling capabilities. More recently, transformer-based models, such as Informer (Zhou et al., 2021) and FedFormer (Zhou et al., 2022), improved long-range dependency modeling for time series data. However, these models tend to be “case-specific,” requiring task-specific training. In contrast, the lastest pre-trained models for time-series data, like MOMENT (Goswami et al., 2024), Timer (Liu et al., 2024b), and TOKEN (Talukder et al., 2024), offer more generalized and adaptable solutions for time-series tasks.
@@ A. Macroeconomic Event Summary
FOMC Html Quarterly 1993.3 $\sim$2024.6 255 2,550
www.federal
reserve.gov
Unemployment
Insurance Claims
PDF, Txt Weekly 2002.10 $\sim$2024.6 913 9,130 oui.doleta.gov
@@ B. Time-Series Data Summary
US Treasury Bond
at 1-Month (USGG1M)
Open, Close, High, Low 5 Min 2013.01 $\sim$2024.06 751,443
US Treasury Bond
at 5-Year (USGG5YR)
Open, Close, High, Low 5 Min 2013.01 $\sim$2024.06 734,773
@@ 3. Problem Formulation
Time Series Data is defined as $\mathcal{X}:=\{X_{1},X_{2},\dots,X_{|\mathcal{X}|}\}$, where each $X_{i}$ represents the numerical data at time step $i$. An event $\mathcal{E}_{i}$ is aligned with a time series segment $\mathcal{X}_{i-\tau:i+\tau}$, where $i$ denotes the time of the releasement of the event, and $\tau$ represents the duration of the time-series segment both preceding and succeeding time step $i$, denoted as $\mathcal{X}_{i-\tau:i+\tau}\mapsto\mathcal{E}_{i}]$. This alignment reflects the time series segment leading up to the event ($\mathcal{X}_{i-\tau}$) and the period during which the event is expected to have an effect ($\mathcal{X}_{i+\tau}$).
@@ 4.1. Dataset Acquisition
In addition to the textual data, studies (CIESLAK et al., 2019; Gertler and Horvath, 2018; Masciandaro et al., 2024; Rosa, 2016; Tadle, 2022; Rosa, 2013; LUCCA and MOENCH, 2015; Rosa, 2011) have demonstrated that the largest market impacts are typically observed in major U.S. stock indexes and Treasury bonds. Therefore, we focused on collecting high-frequency trading time-series data at 5 minute interval for key stock indexes, including the S&P 500 (SPX), Dow Industrial (INDU), NASDAQ (NDX), as well as U.S. Treasury Bond at 1-Month (USGG1M) and Treasury Bond at 5-Year (USGG5YR).
@@ 4.2. Counterfactual Events Generation based on LLM
5. CAMEF Architecture
The CAMEF model integrates both textual and time-series information through a structured architecture consisting of a textual encoder, a time-series encoder, and a forecasting decoder, as depicted in Fig. 3. Each component is detailed below.
@@ 5.1. Textual Modality Encoder ($\textbf{CAMEF}_{\text{Textual}}$)
We encode event scripts using RoBERTa (Liu et al., 2019). Given an input script $\mathbf{E}_{i}={w_{1},w_{2},\dots,w_{m}}$, RoBERTa produces contextual token embeddings:
@@
where $\mathbf{W}^{(1)}\in\mathbb{R}^{768\times 1024}$, $\mathbf{W}^{(2)}\in\mathbb{R}^{1024\times 1024}$, and $\mathbf{W}^{(3)}\in\mathbb{R}^{1024\times 768}$. The final encoding vector $\mathbf{E}_{i}$ for the script is computed as the average of the transformed embeddings, $\mathbf{E}_{i}=\frac{1}{m}\sum_{j=1}^{m}\mathbf{e}_{j}$.
@@ 5.2. Time-Series Modality Encoder
where $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{1024\times 1024},\mathbf{W}_{3}\in\mathbb{R}^{1024\times 768}$ are the weight matrices, and $\mathbf{b}_{1},\mathbf{b}_{2}\in\mathbb{R}^{1024},\mathbf{b}_{3}\in\mathbb{R}^{768}$ are the respective biases. Finally, $\mathbf{Z}_{i}\in\mathbb{R}^{1024}$ serves as the refined vector for the time series segment.
@@ 5.3. Feature Fusion and Time Series Decoder
After obtaining the encoded vectors from the textual and time series data, denoted as $\mathbf{E}_{i}$ and $\mathbf{Z}_{i}$, respectively, we concatenate them to form a unified representation:
(7) $\mathbf{E}_{\text{combined}}=\text{Concat}(\mathbf{E}_{i},\mathbf{Z}_{i}),$
which captures both semantic content from macroeconomic texts and temporal patterns from time series inputs. To fuse these modalities, $\mathbf{E}_{\text{combined}}$ is passed through a two-layer feedforward network with GELU activation:
(8) $\mathbf{E}_{\text{fused}}=\mathbf{W}^{(f2)}\cdot\text{GELU}(\mathbf{W}^{(f1)}\cdot\mathbf{E}*{\text{combined}}+\mathbf{b}^{(f1)})+\mathbf{b}^{(f2)},$
where $\mathbf{W}^{(f1)}\in\mathbb{R}^{(2\times 768)\times 1024}$, $\mathbf{W}^{(f2)}\in\mathbb{R}^{1024\times 1024}$, and $\mathbf{b}^{(f1)},\mathbf{b}^{(f2)}\in\mathbb{R}^{1024}$ are the corresponding weight matrices and bias terms. This fusion block enables interaction across modalities and produces a refined joint embedding for downstream decoding.
We then employ GPT-2 (Radford et al., 2019) as the decoder to decode the fused vector by leveraging its effective auto-regressive ability:
@@
where $\mathbf{H}^{(0)}=\mathbf{E}_{\text{fused}}$, and $f_{\text{GPT2\_layer}}^{(l)}$ represents the transformation function of the $l$-th GPT-2 layer, where $l=12$. After the final layer, the output is normalized using a layer normalization function:
@@ 5.4. Time-Series Forecasting Post-Regressor and Learning Objectives
where $\mathbf{R}^{(0)}=\mathbf{H}_{\text{final}}$, $\mathbf{W}^{(k)}$ and $\mathbf{b}^{(k)}$ are the weight matrix and bias of the $k$-th linear layer, respectively. $k=4$ is the total number of layers in the regressor. The final linear layer maps the representation to a vector of shape $(d\times\text{pred\_len})$, where $d$ is the forecast dimensionality and pred_len is the number of predicted time steps:
@@
Model / Datasets
Forecasting
Length
SP500 (SPX) Dow Industrial (INDU) NASDAQ (NDX) USGG1M USGG5YR
@@ 6.1.3. Test Settings
Textual Causal
Feature
Fusion
GPT2
Decoder
Post-
Regressor
35 70 140 35 70 140 35 70 140 35 70 140 35 70 140 35 70 140
@@ 6.1.4. Implementation Overview
6.2. Experimental Results for Event-Driven Time-Series Forecasting (RQ1)
@@
6.3. Ablation Studies on Model Components (RQ2)
@@ 6.4. Ablation Studies on Different Type of Events (RQ3)
6.5. Parameter Sensitivity Analysis
(a) Effect of counterfactual sample size ($\alpha$)
(b) Effect of post-regressor depth ($k$)
Figure 4. Sensitivity analysis of CAMEF on the SP500 dataset with respect to key hyperparameters.
We examine the impact of two key hyperparameters in CAMEF on SPX dataset with 35 predictive length: the number of counterfactual samples ($\alpha$) and the depth of the post-regressor network ($k$). The number of GPT-2 layers ($l$) remains fixed, as we adopt the pre-trained model without modification. As shown in Figure 4, increasing $\alpha$ enhances forecasting accuracy through stronger contrastive supervision, with performance gains saturating around $\alpha=10$. Similarly, increasing $k$ improves decoding capacity, though with diminishing returns beyond $k=4$. These trends suggest that CAMEF performs robustly across a range of configurations, benefiting from moderate complexity increases without overfitting.